OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 6x = 13

OpenStudy (hba):

ADD 3 ON BOTH SIDES

OpenStudy (anonymous):

what? o_o...

OpenStudy (stephstar101):

It's actually 9 >.< (b/2)^2

OpenStudy (anonymous):

\[a^2 + 2ab + b^2 = (a+b)^2\] 2ab here is 6x, with a = x. b = 3 so b^2 = 9

OpenStudy (anonymous):

(x+3)^2